# A Distributional Treatment of Real2Sim2Real for Vision-Driven Deformable Linear Object Manipulation

### v1

@@ Abstract
We present an integrated (or end-to-end) framework for the Real2Sim2Real problem of manipulating deformable linear objects (DLOs) based on visual perception. Working with a parameterised set of DLOs, we use likelihood-free inference (LFI) to compute the posterior distributions for the physical parameters using which we can approximately simulate the behaviour of each specific DLO. We use these posteriors for domain randomisation while training, in simulation, object-specific visuomotor policies for a visuomotor DLO reaching task, using model-free reinforcement learning. We demonstrate the utility of this approach by deploying sim-trained DLO manipulation policies in the real world in a zero-shot manner, i.e. without any further fine-tuning. In this context, we evaluate the capacity of a prominent LFI method to perform fine classification over the parametric set of DLOs, using only visual and proprioceptive data obtained in a dynamic manipulation trajectory. We then study the implications of the resulting domain distributions in sim-based policy learning and real-world performance.
@@ I Introduction
Deformable linear object (DLO) manipulation is an active area of robotics research that deals with a variety of challenging tasks, such as needle threading, lace tying, cable rearrangement, and lasso throwing [1, 2, 3]. Although the geometry of a DLO is well understood, humans show dexterity in quickly estimating physical parameters, such as length and stiffness, from relatively few visual observations of object manipulation trajectories. This is often used to condition control policies within these estimations [4, 5], enabling efficient adaptive control. Achieving a similar level of robotic dexterity requires dealing with the inherent high dimensionality and nonlinearity of such tasks, challenges that are further exacerbated by the inherent noise in visual servoing [6, 7].
@@
Learning policies in simulation and then performing a Sim2Real [8, 3] deployment follows the hypothesis that the numerous simulated iterations of a given task will robustify the learnt policy to parametric variations, as mentioned above [9]. However, this approach requires one to overcome the “reality gap” of robotic simulators, which is particularly problematic for soft objects.
In this work, we first aim to achieve a reliable Real2Sim [10, 11] calibration of our simulator’s parameters $\boldsymbol{\theta}$ to the real-world object parameters. Then, aiming at robust deployment, we account for any uncertainty about these parameters by exposing our learning algorithm to different hypotheses of $\boldsymbol{\theta}$.
Likelihood-free inference (LFI) deals with solving the inverse problem of probabilistically mapping real-world sensor observations to the respective simulation parameters $\boldsymbol{\theta}$ that are most likely to account for the observations.
Figure 1: Overview of our Real2Sim2Real framework. We perform inference for the posterior distribution $\hat{p}$ over system parameters $\boldsymbol{\theta}$ (Real2Sim). We use $\hat{p}$ to perform domain randomisation while training a PPO agent to perform a visuomotor DLO control task. We deploy and evaluate our sim-trained policy in the real world (Sim2Real).
@@
This involves inferring the multimodal distribution from which we can sample sets of simulation parameters with a high probability of replicating the modelled real-world phenomenon [12]. Any uncertainty over $\boldsymbol{\theta}$ can be modelled as the variances associated with the modes.
@@
This enables us to perform Domain Randomisation (DR), which aims to create a variety of simulated environments, each with randomised system parameters, and then to train a policy that works in all of them for a given control task objective. Assuming that the parameters of the real system is a sample in the distribution associated with the variations seen at training time, it is expected that a policy learnt in simulation under a DR regime will adapt to the dynamics of the real world environment.
@@
Recent literature has demonstrated how distributional embeddings, such as reproducing kernel Hilbert spaces [13], of inferred keypoint trajectories combine low dimensionality and robustness to visual data noise [14], thus enabling robust Real2Sim calibration. On the Sim2Real side, we have seen fruitful model-based Reinforcement Learning (RL) approaches [15, 16] to train deformable object manipulation policies in simulation and deploy them in the real world. However, we have yet to see an end-to-end Real2Sim2Real system which combines the expressiveness of Bayesian inference with the flexibility of model-free RL [17].
We propose an integrated framework addressing a number of different technical considerations (Fig. 1, Alg. 1). We use BayesSim [12] with distributional state embeddings to infer the posteriors over physical parameters $\boldsymbol{\theta}$ of a set of visually observed (real) DLOs of varied length and stiffness. We use each inferred posterior $\hat{p}(\boldsymbol{\theta})$ to perform domain randomisation and train a Sim2Real transferable model-free RL policy to perform a visuomotor “reaching” task with a DLO. We deploy our sim-trained policies in the real world and compare their performance across our set of DLOs.
@@
    We present an end-to-end Real2Sim2Real framework for robust vision-based DLO manipulation.
@@
    We examine the capacity of BayesSim to finely classify the different physical properties of a deformable object drawn from a parametric set.
@@
    We study the implications of different randomisation domains for RL policy learning in simulation and how this translates to real world performance.
Our experiments show that for a parameterised set of DLOs, an integrated distributional treatment of parameter inference, policy training, and zero-shot deployment enables inferring fine differences in physical properties and adapting RL agent behaviour to them.
@@ II-B Domain Randomisation
The goal of RL is to maximise the expected sum of future discounted rewards (by $\gamma$) following a policy $\pi_{\boldsymbol{\beta}}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$, parameterised by $\boldsymbol{\beta}$. We formulate DR over the RL problem as maximising the joint objective:
@@
where $\boldsymbol{\theta}\sim\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$ with respect to policy’s $\boldsymbol{\beta}$. In the inner RL objective formulation, we follow the standard textbook denotations [22].
@@ II-C Kernel Mean Embeddings & Neural Approximation
With kernel trick we avoid operating on infinite-dimensional implicit maps $\phi(\mathbf{x})$, and instead operate on a finite-dimensional Gram matrix $K$: $K_{ij}~{}=~{}k(\mathbf{x}_{i},\mathbf{x}_{j}),\;i,j~{}=~{}1...N$ [23], which, however, is computationally expensive for large datasets ($\mathcal{O}(n^{2})$).
Following [12] (Eq. 12-15) and [14] (Eq. 6) derivations, we can use Random Fourier Features (RFF) [25] to approximate a shift-invariant kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ by a dot product $\hat{\phi}(\mathbf{x})^{T}\hat{\phi}(\mathbf{x}^{\prime})$, which is a more scalable approximation of $K$. $\hat{\phi}(\mathbf{x})$ is a finite-dimensional approximation of $\phi(\mathbf{x})$, with randomised features of the form $\langle\cos(\sigma^{-1}~{}\circ~{}\boldsymbol{\omega}_{m}^{T}\mathbf{x}),\sin(\sigma^{-1}~{}\circ~{}\boldsymbol{\omega}_{m}^{T}\mathbf{x})\rangle$ (cos-sin) or $\cos(\boldsymbol{\omega}_{m}^{T}\mathbf{x}~{}+~{}\mathbf{b}_{m})$ (cos-only) [26].
@@ III Method
III-A Keypoint detection from segmentation images
Our perception module receives an RGB image observation as input, in which it detects the environment objects of interest, i.e. the controlled DLO and the visuomotor control target. We use detection masks to create segmentation images, through which we track our task’s features for parameter inference and policy learning.
We use segmentation images to efficiently learn an unsupervised keypoint detection model, using the transporter method [27], as implemented by [28]. The method’s combination of keypoint-based bottleneck layer and downstream reconstruction task maintains the temporal consistency of the extracted set of keypoints. However, keypoints are inferred in every timeframe, which in practice, particularly when working with real-world image data of deformable objects, creates problems such as keypoint position noise and permutations [14].
III-B Real2Sim with likelihood-free inference
Our problem is that of manipulating a DLO for a visuomotor reaching task, by learning a policy in simulation and deploying in the real-world, without any further fine-tuning. The Real2Sim part of our work deals with calibrating physical parameters that induce the deformable object’s behaviour, such as its dimensions and material behaviour, which are difficult to tune manually. Details such as camera and workspace object placement, controller stiffness parameters, etc. are beyond the scope of our work, although they can be incorporated as additional tunable parameters in extensions of our method.
Following Algorithm 1, we begin by assuming a uniform proposal prior $\tilde{p}(\boldsymbol{\theta})$, which we use to initialise our reference prior $p_{0}=\tilde{p}$. We then perform domain randomisation as $\boldsymbol{\theta}\sim p_{0}$, while training in simulation an initial policy $\pi_{\boldsymbol{\beta}_{0}}$ for our task. We execute a rollout of $\pi_{\boldsymbol{\beta}_{0}}$ in the real world environment, to collect a real trajectory $\mathbf{x}^{r}$ while manipulating a specific DLO. We perform LFI through multiple BayesSim iterations [21, 14]. In each inference iteration $i$, we use $\pi_{\boldsymbol{\beta}_{0}}$ as a data collection policy, running $N$ rollouts of $\pi_{\boldsymbol{\beta}_{0}}$ in simulation to collect a dataset $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N},\boldsymbol{\theta}\sim\tilde{p}$, which we use to train our BayesSim conditional density function $q_{\phi}$. We use our $q_{\phi}$ and $\mathbf{x}^{r}$ to compute the posterior $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$. We then update our reference prior $p_{i}=\hat{p}$ and loop again.
Figure 2: An overview of our policy rollout and trajectory perception method, with keypoint extraction from segmentation images, and keypoint trajectory cos-only kernel mean embeddings using the RKHS-net layer (BayesSim-RKHS), illustration inspired by [14].
III-C Keypoint trajectories in RKHS
The input of the RKHS-net layer is a vector of samples of a distribution. In this work, we compute the distributional embedding for a noisy trajectory of keypoints $\mathbf{x}~{}=~{}(\mathbf{x_{1}},...,\mathbf{x_{n}})$, where each keypoint is defined in a $2$D RGB pixel space.
III-D Policy Learning and Sim2Real Deployment
We use Proximal Policy Optimisation (PPO) [17] as our reference model-free on-policy RL algorithm. PPO is designed to maximise the expected reward with a clipped surrogate objective to prevent large updates. It samples actions from a Gaussian distribution, which is parameterised during policy learning. We perform policy learning in simulation, performing domain randomisation using our likelihood-free inference reference prior $p$ as a task parameterisation sampler, where $p$ is either the default uniform distribution $\mathit{U}$, or an inferred MoG posterior (Alg. 1, line 12).
@@ III-D Policy Learning and Sim2Real Deployment
6:// 1. Real2Sim DLO parameter inference
@@ IV Experimental Setup
    How does this impact translate into object-centric agent performance during real-world deployment?
@@ IV-A Reference Task
To evaluate the parameter inference performance of our set-up, as well as the implications of the inferred parameter posterior $\hat{p}(\boldsymbol{\theta})$ to policy learning and deployment, we design a visuomotor task which studies the dynamic manipulation of a DLO towards a target in its workspace. Our task is inspired by established dynamic rope manipulation tasks [29, 2].
We set up our task with a robot arm that picks up the DLO from a designated position on the table $\mathbf{x}_{0}$ by grasping it near one of its tips, and raises it to a designated height $h_{0}$. Both $\mathbf{x}_{0}$ and $h_{0}$, as well as our initial DLO pose, remain fixed for all of our simulation and real experiment iterations. The execution of our policy begins immediately after the object has been picked up and raised to $h_{0}$. Thus, the dynamic nature of our task is constituted by the underactuated manner in which we control our DLO, having it dangling from one of its tips, as well as the momentum forces acting on its body ever since it was raised from the table.
The positions of both the DLO and the target are tracked in the 2D pixel space of an RGB image. More specifically, the DLO is tracked by an inferred cluster of $4$ keypoints, extracted over a segmentation image, and the target with the computed pixel-space median of its own segmentation mask, which we treat as an artificial $5$th keypoint. In both our parameter inference (real2sim) and policy training and deployment (sim2real) experiments, we position-control our Panda end-effector (EEF) by navigating it to a target position through inverse kinematics. Our simulator uses the IsaacGym attractors implementation, whereas in the real world we use a Cartesian impedance controller [30].

### v2

# A Distributional Treatment of Real2Sim2Real for Object-Centric Agent Adaptation in Vision-Driven Deformable Linear Object Manipulation

@@ Abstract
We present an integrated (or end-to-end) framework for the Real2Sim2Real problem of manipulating deformable linear objects (DLOs) based on visual perception. Working with a parameterised set of DLOs, we use likelihood-free inference (LFI) to compute the posterior distributions for the physical parameters using which we can approximately simulate the behaviour of each specific DLO. We use these posteriors for domain randomisation while training, in simulation, object-specific visuomotor policies (i.e. assuming only visual and proprioceptive sensory) for a DLO reaching task, using model-free reinforcement learning. We demonstrate the utility of this approach by deploying sim-trained DLO manipulation policies in the real world in a zero-shot manner, i.e. without any further fine-tuning. In this context, we evaluate the capacity of a prominent LFI method to perform fine classification over the parametric set of DLOs, using only visual and proprioceptive data obtained in a dynamic manipulation trajectory. We then study the implications of the resulting domain distributions in sim-based policy learning and real-world performance.
@@ I Introduction
Deformable linear object (DLO) manipulation is an active area of robotics research that deals with challenging tasks, such as lace tying, surgical suturing, and rope whipping [1, 2, 3]. Even when we know a DLO geometry, adapting a policy to its finer physical parameters is crucial to achieve precise physical interactions. For this paper, let us consider a visuomotor reaching task that seemingly requires less dexterity, but indicates readiness to scale up toward the aforementioned tasks. We hold a DLO above a table and want to guide its entire body toward a 2D visual target within a fixed time horizon, using visual and proprioceptive observations. Our objective can be quantified as minimising the overall distance of the DLO body to the target (Fig. 1).
Let us think of how efficiently we would solve such a task; intuitively estimating DLO length and stiffness, from a few visual observations of manipulation trajectories, and adapting our actions to this estimation. An effective sequence of such actions would maximise whole-body convergence to the target for a given DLO parameterisation, while implicitly minimising adversarial factors, such as DLO drag on the table surface. Achieving a similar level of robotic dexterity requires conditioning control policies to physical parameter estimations [4, 5], while dealing with the high dimensionality and nonlinearity of DLO states, challenges further exacerbated by the inherent noise in visual servoing [6, 7].
Figure 1: Overview of our Real2Sim2Real framework (top). We perform LFI for the posterior distribution $\hat{p}$ over system parameters (Real2Sim). We use $\hat{p}$ to perform domain randomisation while training a PPO agent to perform a DLO reaching task. We deploy and evaluate our sim-trained policy in the real world (Sim2Real). We experiment with a DLO reaching task, which is an abstraction of more intricate physical interactions, such as whipping the top cube of a stack (left). We demonstrate strong object-centric agent adaptation relative to the underlying domain distribution $\hat{p}$. For example (bottom), an effective task policy should not drag the DLO on the table (a), but reach for the green target (cross) with a larger part of its body (b).
@@
Learning policies in simulation and then performing a Sim2Real deployment [8, 3] follows the hypothesis that multiple simulated iterations of a task will robustify the learnt policy to parametric variations [9]. This requires overcoming the “reality gap” of robotic simulators, which is particularly problematic for soft objects. Thus, we first need a reliable Real2Sim calibration [10, 11] of the simulator parameters $\boldsymbol{\theta}$ to real-world parameters. Then, for robust deployment, we account for any uncertainty about these parameters by exposing our learning algorithm to different $\boldsymbol{\theta}$ hypotheses.
@@
Likelihood-free inference (LFI) deals with solving the inverse problem of probabilistically mapping real-world sensor observations to the respective simulation parameters $\boldsymbol{\theta}$ that are most likely to account for the observations. This involves inferring the multimodal distribution $\hat{p}(\boldsymbol{\theta})$ from which we can sample sets of simulation parameters with a high probability of replicating the modelled real-world phenomenon [12]. Any uncertainty over $\boldsymbol{\theta}$ can be modelled as the variances associated with the modes.
@@
This enables us to use $\hat{p}(\boldsymbol{\theta})$ as a sampler while performing Domain Randomisation (DR). DR aims to create a variety of simulated environments, each with randomised system parameters, and then to train a policy that works in all of them for a given control task objective. Assuming that the parameters of the real system is a sample in the distribution associated with the variations seen at training time, i.e. $\hat{p}(\boldsymbol{\theta})$, we can expect that a policy learnt in simulation under this regime will adapt to the real-world dynamics.
@@
Recent literature has demonstrated how distributional embeddings, such as reproducing kernel Hilbert spaces [13] of inferred low-dimensional keypoint trajectories, provide robustness to visual data noise and permutation invariance [14], thus enabling robust Real2Sim calibration. On the Sim2Real side, we have seen fruitful model-based Reinforcement Learning (RL) approaches [15, 16] to train deformable object manipulation policies in simulation and deploy them in the real world. However, we have yet to see an end-to-end Real2Sim2Real system which combines the expressiveness of Bayesian inference with the flexibility of model-free RL [17].
@@
    We propose an end-to-end Real2Sim2Real framework for robust vision-based DLO manipulation.
@@
    We examine the capacity of BayesSim [12] with distributional state embeddings to finely classify the physical properties of a DLO drawn from a parametric set.
@@
    We study the implications of different randomisation domains for model-free RL policy learning in simulation and how this translates to real-world performance.
Our experiments show that for a parameterised set of (real) DLOs, an integrated distributional treatment of parameter inference, policy training, and zero-shot deployment enables inferring fine differences in physical properties and adapting RL agent behaviour to them.
@@ II-B Domain Randomisation
The goal of RL is to maximise the expected sum of future discounted rewards (by $\gamma$) following a policy $\pi_{\boldsymbol{\beta}}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$, parameterised by $\boldsymbol{\beta}$ and sampling action $\mathbf{a}_{t}$ given state $\mathbf{s}_{t}$. We formulate DR over the RL problem as maximising the joint objective:
@@
with respect to policy’s $\boldsymbol{\beta}$, where $\boldsymbol{\theta}\sim\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$ and $\boldsymbol{\eta}=\{\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{o}_{t}\}_{t=0}^{T-1}$ the history of observation $\mathbf{o}_{t}$, action $\mathbf{a}_{t}$ pairs over time horizon $T$. In this paper, we consider $\mathbf{s}_{t}\equiv\mathbf{o}_{t}$.
@@ II-C Kernel Mean Embeddings & Neural Approximation
With kernel trick we avoid operating on infinite-dimensional implicit maps $\phi(\mathbf{x})$, and instead operate on a finite-dimensional Gram matrix $K_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j}),i,j=1...N$ [22], which, however, is computationally expensive for large datasets ($\mathcal{O}(n^{2})$).
Following [12] (Eq. 12-15) and [14] (Eq. 6) derivations, we can use Random Fourier Features (RFF) [24] to approximate a shift-invariant kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ by a dot product $\hat{\phi}(\mathbf{x})^{T}\hat{\phi}(\mathbf{x}^{\prime})$, which is a more scalable approximation of $K$. $\hat{\phi}(\mathbf{x})$ is a finite-dimensional approximation of $\phi(\mathbf{x})$, with randomised features of the form $\cos(\boldsymbol{\omega}_{m}^{T}\mathbf{x}+\mathbf{b}_{m})$ (cos-only) or $\langle\cos(\sigma^{-1}\circ\boldsymbol{\omega}_{m}^{T}\mathbf{x}),\sin(\sigma^{-1}\circ\boldsymbol{\omega}_{m}^{T}\mathbf{x})\rangle$ (cos-sin) [25].
@@ III Method
6:// 1. Real2Sim DLO parameter inference (LFI)
@@
III-A Keypoint detection from segmentation images
Our perception module receives an RGB image as input, in which it detects the environment objects of interest, i.e. the controlled DLO and the task’s target. We use detection masks to create segmentation images, through which we track our task’s features for parameter inference and policy learning.
We use segmentation images to efficiently learn an unsupervised keypoint detection model, using the transporter method [26], as implemented by [27]. The resulting keypoints, although temporally consistent, are inferred in every timeframe, creating problems such as pixel position noise and permutations [14], particularly when working with real-world image data of deformable objects. The lack of permutation invariance hinders the creation of reward functions specific to DLO parts, such as the tip [28, 2]. Focussing on whole-body guidance abstracts this challenge and raises the importance of inferred posteriors to implicitly learn successful behaviours using less elaborate reward functions.
III-B Real2Sim with likelihood-free inference
Considering our prefaced DLO visuomotor reaching task, we want to learn a policy in simulation and deploy it in the real world, without any further fine-tuning. The Real2Sim part of our work deals with calibrating physical parameters that induce the deformable object’s behaviour, such as its dimensions and material behaviour, which are difficult to tune manually. Details such as camera and workspace object placement, controller stiffness parameters, etc. are beyond the scope of our work, although they can be incorporated as additional tunable parameters in extensions of our method.
Following Algorithm 1, we begin by assuming a uniform proposal prior $\tilde{p}(\boldsymbol{\theta})$, which we use to initialise our reference prior $p_{0}=\tilde{p}$. We then perform domain randomisation as $\boldsymbol{\theta}\sim p_{0}$, while training in simulation an initial policy $\pi_{\boldsymbol{\beta}_{0}}$ for our task. We execute a rollout of $\pi_{\boldsymbol{\beta}_{0}}$ in the real world environment, to collect a real trajectory $\mathbf{x}^{r}$ while manipulating a specific DLO. We perform LFI through multiple BayesSim iterations [21, 14]. In each inference iteration $i$, we use $\pi_{\boldsymbol{\beta}_{0}}$ as a data collection policy, running $N$ rollouts of $\pi_{\boldsymbol{\beta}_{0}}$ in simulation to collect a dataset $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N},\boldsymbol{\theta}\sim\tilde{p}$, on which we train our BayesSim conditional density function $q_{\phi}$. We use our $q_{\phi}$ and $\mathbf{x}^{r}$ to compute the posterior $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$. We then update our reference prior $p_{i}=\hat{p}$ and loop again.
Figure 2: Overview of our policy rollout and trajectory perception method. In each policy step we extract keypoints from segmentation images. Once all rollouts are completed, for each keypoint trajectory we compute the cos-only kernel mean embeddings of each step’s keypoints using the RKHS-net layer (BayesSim-RKHS). Margins colour coding indicates association between sample images and algorithm parts. Illustration inspired by [14].
III-C Keypoint trajectories in RKHS
The input of the RKHS-net layer is a vector of samples of a distribution. In this work, we compute the distributional embedding of a noisy keypoint trajectory $\mathbf{x}=(\mathbf{x_{1}},...,\mathbf{x_{n}})$, each keypoint defined in a $2$D RGB pixel space.
III-D Policy Learning and Sim2Real Deployment
We use Proximal Policy Optimisation (PPO) [17] as our reference model-free on-policy RL algorithm. PPO maximises the expected reward with a clipped surrogate objective to prevent large updates. It samples actions from a Gaussian distribution, which is parameterised during policy learning. We learn policies in simulation, performing domain randomisation using our LFI reference prior $p$ as a task parameterisation sampler, where $p$ is either the default uniform distribution $\mathit{U}$, or an inferred MoG posterior (Alg. 1, l. 12).
@@ IV Experimental Setup
    How does this translate into object-centric agent performance adaptation during real-world deployment?
@@ IV-A Reference Task
To evaluate the parameter inference performance of our setup, and the implications of the inferred parameter posterior $\hat{p}(\boldsymbol{\theta})$ for policy learning and deployment, we design a visuomotor reaching task that studies the guidance of the entire DLO body towards a visual target.
We set up our task with a robot arm that picks up the DLO from a designated position on the table $\mathbf{x}_{0}$ by grasping it near one of its tips, and raises it to a designated height $h_{0}$. Both $\mathbf{x}_{0}$ and $h_{0}$, as well as our initial DLO pose, remain fixed for all of our simulation and real experiment iterations. The execution of our policy begins immediately after the object has been picked up and raised to $h_{0}$. Thus, the dynamic nature of our task is constituted by the underactuated manner in which we control our DLO, which dangles from one of its tips, as well as the momentum forces (drag, inertia) acting on its body ever since it was raised from the table.
The positions of the DLO and the target are tracked in the 2D pixel space of an RGB image. The DLO is tracked by an inferred cluster of $4$ keypoints, extracted over a segmentation image, and the target with the computed pixel space median of its own segmentation mask, which we treat as an artificial $5$th keypoint. In both our parameter inference (real2sim) and policy training and deployment (sim2real) experiments, we position-control our Panda end-effector (EEF) by commanding its Cartesian pose. Our simulator uses the IsaacGym attractors implementation, whereas in the real world we use a Cartesian impedance controller [29].